If 2 body has same kinetic energy . what is the ratio of its momemtum?

Grade:12

1 Answers

Eshan
5 years ago
Dear student,

Kinetic energy of a body=\dfrac{1}{2}mv^2=\dfrac{(mv)^2}{2m}=\dfrac{p^2}{2m}

Hence for equal kinetic energy,p\propto \sqrt{m}

\implies \dfrac{p_2}{p_1}= \dfrac{\sqrt{m_2}}{\sqrt{m_1}}